OpenStudy (kaiya):

Find the derivative of the function. r(x) = (e^6x-1)^8 I am having a hard time understanding derivatives in my Calculus class right now. I would very much appreciate if someone could help me solve this problem. It's one of the ones I'm stuck on at the moment. =(

OpenStudy (anonymous):

chain. rule.

OpenStudy (anonymous):

you should first use the one of the Chain Rule--General Power Rule: d/dx [u^n]=nu^(n-1)u' set (e^6x-1) as u, so you will get: r(x)=u^8 then you apply the power rule to it, you will get: r'(x)=8u^7u' so what is u' ? u'=d/dx [e^6x-1] the derivative of e^(6x)-1 is 6e^(6x) because d/dx [e^u] = u' e^u now you can get the answer: r'(x)=8(e^6x-1)^7(6e^(6x))

OpenStudy (kaiya):

wow ty so much cc, your answer was so very helpful, and explained it all perfectly!! ty<3

OpenStudy (anonymous):

you are welcome~ enjoy your calculus :3